Globalization and mythology

R.W. Oldford University of Waterloo

May 9, 2002

“This is a time of great opportunity. What some call globalization is in fact the triumph of human liberty across national borders. We have today the chance to prove that freedom can work not just in the new world or old world, but in the whole world. Our great challenge is to include all the world’s poor in an expanding circle of development throughout all the Americas and all of Asia and all of Africa. Such a world will enjoy greater freedom and prosperity, and is far more likely to be at peace.”

George W. Bush, President of the United States, delivered these words July 21, 2001 in his weekly radio address to the nation. Uttered by the most powerful person in the world, they command our attention. But more than that, they actually merit our attention.

The reasons are several, but primary amongst them must be the language itself. The President’s speechwriters have produced a minor masterpiece of rhetoric.

Liberty, freedom, prosperity, peace. Who could oppose any of these? They are core to any democracy and do indeed transcend national boundaries. We all look forward to the triumphal march of “human liberty across national borders”. How reassuring that this turns out to be globalization.

There is the ring of historical inevitability here. Freedom has been proven to work in the new world; it has been proven to work in the old world. Where first, it is not made clear, nor does it need to be – the American audience will decide. What is made clear is that freedom can be proven to work in the whole world. In the end, freedom will win out.

It might not happen now, although the time is opportune. It is not certain, but we do have a chance. It won’t be easy but, if we are up to the great challenge, the result will be the inclusion of “all the world’s poor in an expanding circle of development throughout all the Americas” (i.e. South and Central America) “and all of Asia and all of Africa” (i.e. there are no poor in North America, Europe, or Australasia). If we hesitate, the opportunity could slip away. It is the good fight, fit for a hero.

Consider the backdrop to this speech. The U.S. President is on foreign soil, meeting with foreign heads of state, his debut on the international stage. On the streets outside mobs of tens of thousands confront the democratically elected leaders inside; they are demanding an end to globalization. There is violence. Against these odds, clear on the justness of his cause for freedom, the President appeals directly to the “American people” telling it like it is. Imagine James Stewart or Gregory Peck in the role or, nowadays, Denzel Washington.

The brilliance of the speech lies in the fact that nowhere, not in the segment I have quoted nor anywhere in its full text, is there given an argument for globalization. Instead, globalization is framed in a setting and conjoined with dearly held human values, both of which figure prominently in the American national mythology. All would resonate with listeners back home. They know the story; they know the outcome.

Nature and Role of Mythology

Bruce Meyer (2000, p.14), writing on the great books of Western Civilization, has said:

“Myths are rarely logical, yet they seem to say a great deal about the way our minds work. And when we cannot come up with fact or scientific laws to explain a person, an event, or why we do things in a certain way, our imaginations take over and create an answer for us. … Our minds desperately want to make sense of what we see and what we experience, but when logic, science, and fact fall short or simply aren’t available, the mind goes to work and invents an appreciable alternative. That is what mythology is: information which fills the vacuum of fact.”

I think that George W. Bush believes what he says about globalization. Mythology, firmly engrained, will draw those facts to it which reinforce it while ignoring others. Francis Bacon pointed this out nearly 400 years ago when he wrote

“The human understanding, once it has adopted opinions, either because they were already accepted and believed, or because it likes them, draws everything else to support and agree with them.” … Aphorism 46, Novum Organum, 1620.

Mythology does not belong to an individual, to exist it must be shared. When the prevailing mythology is consistent with a particular viewpoint, there is little need to find facts which support that viewpoint and no motivation to find facts to the contrary.

The first great book in Bruce Meyer’s canon is, not surprisingly, the Bible. Meyer points out that

“The Bible reflects a number of different approaches to the definition of mythology, and all of them appeal to our imaginations, perhaps because they reflect the way our imaginations work.” Meyer (ibid, p. 14)

He goes on to describe explanatory mythology whereby the structure of the universe is described with God firmly in charge and order taking precedence over chaos. The explanatory mythology describes why things are as they are. It seldom describes whether things could have been different; there are no alternatives.

Next is instructional mythology, which provides the guidelines to be followed, the “how to” for readers to follow so as to be in harmony with the explanatory mythology, that is with God’s plan.

The ability to follow guidelines is a fundamental building block to the third type of mythology, the legal or moral mythology which lays out the dos and don’ts of human behaviour. Legal mythology determines the boundaries of acceptable conduct.

Finally, there is historical mythology, the “chronicle of process” as Meyer puts it. The Bible describes the history of one group of people entirely from their perspective; no other group or perspective matters. The narrative selectively expands on different time periods, telling in varying detail the ongoing relationship between God and his people. The reader recognizes himself in the story, that his life is “a reflection in miniature of the grand scale.”

No book has had greater impact on Western Civilization. Every home would endeavour to possess at least this one book. Individuals would see themselves within the mythology and with repetition the stories would resonate in society.

For nearly two millennia, the Bible formed the fundamental mythology common to Western Civilization. The Holy Roman Empire, the Crusades of the Middle Ages, and the Age of European Empires, have all been able to turn to this mythology for comfort, support and justification. In other times and other places, other mythologies have served as well. One need only think of the Cold War to bring us to modern times and the role played by conflicting mythologies.

When unspeakable acts^{1} are practised, or merely witnessed, both the individual and the group gain mutual reassurance from their understanding of the larger picture, of the plan for the world, and of one’s place in it. In the moment, the framework helps explain the difficulties away and, as time passes, to forget them altogether. And so historical mythology will trump history.

If nothing else, our explanatory and historical mythologies offer the comfort that we have brought civilization to another, unenlightened, part of the world. In this light, one begins to appreciate the value of the conjunction of development and globalization in President Bush’s speech.

Nation States and Corporations

For most of history, the groups which go global have been nation states. Trade is one way to do this but subjugation seems historically to have been preferred. Nation states are rightly concerned about their own security and will form strategic alliances so as to balance and maintain power.

1Consider one example, that of Francisco Pizarro and his conquest of the Incan civilization for Spain.

In 1532, the Incas were experiencing a civil war as two brothers fought for the crown. King Atahualpa was already upset with the behaviour of Pizarro and his men in Peru when Pizarro arranged a meeting with him in the main square of the town of Caxamalca.

When Atahualpa arrived with an unarmed entourage of thousands he was met only by a Dominican friar, with Bible in hand, and an interpreter by his side. The Dominican explained the nature and superiority of the Christian faith to Atahualpa, starting at Creation and working his way through to the doctrine of the Trinity. The conclusion was that the Incas were to adopt the faith and be subjugated by Spain and its king.

Atahualpa examined the Bible and threw it to the ground. This was the signal event for Pizarro and his men, who hidden from view had surrounded the square, to attack with horse, sword and cannon slaughtering the unarmed thousands and capturing Atahualpa. On the spot, the priest absolved Pizarro and his men for the acts they were about to perform.

There are two remarkable footnotes to this story. First, the information is that actually reported in the Spanish records of the time. There seems to have been little reluctance in recording such events. Second, an internet search for “Pizarro” and “Bible” turned up a recent sermon entitled “By faith, not by sight”. Remarkably, the pastor chose to clarify what was implied by “Pick up your cross, deny yourself, and follow Me” by relating how Pizarro’s men had to choose between enduring continued hardship to follow him on to Peru, or to turn back. The sermon quotes an unnamed ‘historical’ source: “A boat turned immediately toward home with the quitters, and Pizarro and thirteen men turned the other ship to Peru, and changed the course of history.” The pastor would have us follow Pizarro’s example, historical mythology trumping history.

Conversely, if easy pickings are to be had, they will take them.

To hold power over existing states, Machiavelli put it plainly

“Whenever those states which have been acquired … have been accustomed to live under their own laws and in freedom, there are three courses for those who wish to hold them: the first is to ruin them, the next is to reside there in person, the third is to permit them to live under their own laws, drawing a tribute, and establishing within it an oligarchy which will keep it friendly to you. Because such a government, being created by the prince, knows that it cannot stand without his friendship and interest, and does its utmost to support him; and therefore he who would keep a city, accustomed to freedom, will hold it more easily by the means of its own citizens than in any other way.”

Nicolo Machiavelli, from The Prince, Written c. 1505, published 1515; Translated by W. K.

Marriott

Such tributary states are best, for they are low maintenance and the problem of dissent is transferred to the oligarchy in power.

The quietening of dissent is an important concern to those who wish to wield and hold power. Authoritarian states and organizations have access to denial of life or livelihood but as Machiavelli points out, there is a better solution – tributary oligarchic states or organizations. Applicable at any level, this solution works best in conjunction with a sympathetic mythology. Fostering such a mythology is immeasurably helped by control of, or at least influence over, the flow of information. Power dissipates as control over information is lost.

Active for hundreds of years, corporations are no newcomers as groups who seek to globalize their operations. Winston Churchill (1956, p. 103) wrote

“The British Empire in India, which was to be painfully built up in the course of the next three centuries, owes its origins to the charter granted by Queen Elizabeth to a group of London merchants and financiers in the year 1600.”

This is the East India company, a company which Churchill describes as “a small and struggling affair … [having] … at the start … a capital of only 72,000 pounds” (ibid). This mythology is seen for what it is a page later where he reports that the annual revenue for the English Government at the time is 300,000 pounds. A similarly “small and struggling” American corporation starting out in 1999 would have had an initial capital of only 456 billion dollars! Promoting the idea that world markets are developed by risk-taking adventurers is disingenuous.

Nation states and corporations have always worked hand in glove. However, it has not always been clear which is the hand and which is the glove. The following quotation has been attributed to Abraham Lincoln in 1864: ^{2} “I see in the near future a crisis approaching that unnerves me and causes me to tremble for the safety of my country … corporations have been enthroned and an era of corruption in high places will follow, and the money of the country will endeavor to prolong its reign by working upon the prejudices of the people until all wealth is aggregated in a few hands and the republic is destroyed. I feel at this moment more anxiety for the safety of my country than ever before, even in the midst of war.”

2Lincoln is an icon of American mythology and much effort has been expended to cultivate and maintain the character of the icon. It is no surprise then that the authenticity of this quote is the subject of some debate. It does not appear, for example, in the Abraham Lincoln Association’s collected works (Basler, 1953) yet does appear in the Lincoln Encyclopedia (Shaw, 1950, p. 40).

For our purposes, it is sufficient that it is a quote that is possibly contemporary with Lincoln. Shibley (1896, p.

282-3) notes that the quote “was published in a collection of Lincoln’s Sayings, years and years ago, by Mr. Jesse Harper. Mr. Harper is still living and is a respected citizen of Danville, Ill.” It is also of some interest that Shibley is arguing against the destabilization of money and consequently of exchange rates. Because the letter has not been physically produced, doubt of its authenticity must remain.

U.S. President Abraham Lincoln, Nov. 21, 1864 (letter to Col. William F. Elkins)

In 1868, with the emancipation of the slaves, the U.S. Constitution was amended to provide equal protection for “[a]ll persons born or naturalized in the United States” and, in particular, protection against differential treatment by state governments. In 1886, the U.S. Supreme Court made a landmark ruling that these same protections applied to a corporation, that within the meaning of its Equal Protection Clause a corporation is a ‘person’. The justices had no interest in argument for this interpretation and presented none — they found it to be self-evident. With some dissenting minority opinion in 1949, the decision has been upheld ever since and the Supreme Court has struck down any state action which conflicted with ‘corporate rights’ under the Fourteenth Amendment.

Legal mythology is not only the set of rules which are explicitly laid out, but also the interpretation of these rules. The mythology lies in the belief that rules can be written which could be, or should be, universally and consistently applied in all contexts – even those which were not imagined by the original authors. In the case of the fourteenth amendment the U.S. legal mythology, designed to serve humans, was easily adapted to serve corporations.

In a similar vein, rules have been developed for nation-states by international organizations such as the World Bank, the International Monetary Fund (IMF), and the World Trade Organization (previously the GATT), which are more in keeping with those for corporations. Not surprisingly, the rules serve corporations well. Moreover the consistency which legal mythology demands means these rules seem able to prevent nation-states from seriously considering bold imaginative steps such as forgiving the debt of the world’s poorest nations.

Similarly, the clamour for corporate deregulation can be seen as adjusting the instructional mythology so as to be geared more toward corporate interests. Reducing taxes and deregulation reduce the power of the government and consequently increase the relative power of the corporations. The public interest seems to lie outside the framework for discussion, somewhere at the margin of power.

Common characteristics

Groups (nation states or corporations) seeking to globalize seem to manifest certain common characteristics.

First amongst these seems to be an existence of a prevailing mythology which is sympathetic to its ends. Which one is the child of the other matters little as the relationship seems to be symbiotic. Prevailing mythology seems to be a natural human crutch.

Second is the notable concentration of power. We have all witnessed the mega-mergers of large corporations in the past decade or so. As with nation states, strategic alliances abound. Corporate boards of directors show a remarkable degree of overlap in their membership. Competition and risk are not as highly valued as stability and predictability. If monopoly is a feasible option it will be taken. If not, the market is controlled by an oligarchy of corporations, each only jostling the other over their fairly stable market shares. The number of shares typically dwarf the number of distinct shareholders; international financial organizations (e.g. IMF) organize their voting in much the same way. Corporate contributions to political campaigns are a large fraction of the money raised. Politicians, when out of power are prominent on the boards of large corporations.

Third, concentrated power is ademocratic, often being undemocratic in its own organization and ambivalent to the political organization of those it deals with. It also leads to a certain detachment with respect to people, organizations, or nations at the margin of power. Downsizing of corporations after merger is one example. Efficiency treated as an end in itself rather than a means to an end is another. Use of child labour in a country with laxer labour laws is yet another.

Fourth, there is value in the quietening of dissent. Control by force, physical or economic, has long been traditional but need not be invoked where a sympathetic prevailing mythology exists and is stable. Whether intentionally, or naturally because it seems self-evident, power will nurture and promulgate a mythology sympathetic to itself. When facts are either absent or complex, this will be particularly effective. And because the mythology must be widely shared, it is important to power to have influence over the information available to the public.

It matters little whether any of these characteristics have come about by conspiracy or as a natural consequence of human and social behaviour. What is important is that societies and individuals be aware of these characteristics so that steps can be taken to identify and avoid their undesirable consequences. Power must be purposely balanced so that it is the society affected which is able to make conscious informed open and democratic decisions.

Concern for the future

Looking to the future, there are many targets for concern but one which is fundamental and universal is the control and communication of information. It is in the rythym and repetition of information that mythologies flourish. The natural inertia of a prevailing mythology makes it challenging to change, let alone expose as mythology. So it is especially worrisome when those in power both benefit from the prevailing mythology and are able to dominate its communication.^{3}

3Consider the following.

The Walt Disney Company is a global media giant with over US $25.4 B in total revenue in 2000. Like other news corporations, Disney’s broadcast news arm, ABC has been trying to have its reporters on the ground in Afghanistan but has met with no success. U.S. Military recently hired Disney producer Jerry Bruckheimer to produce a 13 week prime-time ‘reality TV’ show on the war in Afganistan. Bruckheimer, a former advertising executive, is known for directing or producing such notable box-office hits as ‘Top Gun’, ‘Pearl Harbor’ and more recently ‘Black Hawk Down’ a quite positive portrayal of the American involvement in Somalia. Disney rebuffed the complaints of its ABC news division.

Between jobs as a U.S. General in Desert Storm and George W. Bush’s Secretary of State, Colin Powell served on the board of directors of the internet service provider “America Online”, or AOL. In 2001, AOL petitioned to take over the broadcast and publication giant Time-Warner to create AOL-Time-Warner. To effect the deal, the approval of the Federal Communications Commission (a panel of five politically appointed commisioners) was required. One of the Republican commisioners was no other than Michael Powell, son of Colin Powell. With some complaint from others in the industry, the deal was approved with a single significant condition. Because of AOL’s dominant share of the American internet market, the FCC required AOL to “open up” its proprietary e-mailing software, called ‘Instant Messager’, so that other internet providers could make use of it. Notably Michael Powell dissented. Once Bush became President, Michael Powell was made the Chair of the FCC and the Instant Messenger requirement quickly

Soon there will be a convergence of digital technologies which will deliver traditional content to your home through either the internet via your computer or more likely in the near future through your digital TV. The same network will be able to to serve telephony, text, video and audio. Like the internet, the communication will be two way.

The good news is that this will offer you many more possibilities than are now available and the service can be tailored minute by minute to your own tastes. The bad news is the same as the good news. It means that your behaviour on this device is monitored so that advertising can be directed at you according to how your behaviour fits some profile.

Unlike the introduction of previous mass media technologies (e.g. radio, television) there has been little public debate on how this media might be designed to serve the public interest. This is most worrying because whoever controls the principal entry point to this world of wonder also controls the information which is most easily accessed.

On the internet, this kind of entry point is called a ‘portal’ and is tailored to you and your pattern of use. All internet service providers and computer manufacturers now have portals which come automatically with their service and can be very difficult for the non-technical to override.

Absent any serious public debate, this situation is likely to carry over to two way digital television with its much greater reach and with its small number of ‘vertically integrated’ media giants in charge. Marshall McLuhan’s ‘global village’ is all very well and exciting, but should its town hall be organized by Disney?

Close

Historically, globalization has been of two broad types: deliberate and accidental. Mostly I’ve been been talking about the deliberate kind associated with power. The accidental or incidental type is often an unintended consequence of deliberate globalization.

Examples abound. There is the spread of disease through the movement of humans and animals (from the plague, to small pox, to HIV and BSE), the destruction of species and ecosystems, the dominance and destruction of peoples and cultures, the increased volatility of currency markets, and global environmental changes.

“Globalize” is a transitive verb, moving a thing from local to global circumstances. What constitutes globalization is anything which effects this transition. Three things matter: what is it that is globalized, how is it that the globalization is effected, and what is the effect of the globalization.

Universal opposition or support is not decidable. It depends on the answer to these questions for each particular instance. Time taken in consideration is time well spent. Complexity needs to be embraced and understood without any attendant simplifying mythology. Decisions need to be prudent and power local and democratic.

References

Bacon, F (1620). Novum Organum (Translated and edited by Peter Urbach and John Gibson, 1994) Peru Illinois, Open Court.

dropped.

Basler, R.P., editor (1953). The Collected Works of Abraham Lincoln, VIII 1864-1865. New Brunswick, New Jersey. Rutgers University Press. Bush, G.W. (2001). Radio Address to the Nation, July 21, 2001. http://www.whitehouse.gov/news/ releases/2001/07/20010721.html Churchill, W.S. (1956). A History of the English Speaking Peoples, Vol. II, The New World, New York, Bantam Books. Machiavelli, N. (1515). The Prince. Meyer, B. (2000). The Golden Thread: A Reader’s Journey through the Great Books Toronto, HarperCollins. Shaw, A.H., editor (1950). The Lincoln Encyclopedia. New York. The Macmillan Company. Shibley, G.H. (1896). The Money Question. Chicago, Ill.. The Stable Money Publishing Company. U.S. Constitution: 14th Amendment (1868). Rights, Guaranteed Privileges and Immunities of Citizenship, Due Process and Equal Protection, Section 1. U.S. Supreme Court decision (1886). Santa Clara Co. v. Southern Pacific R. Co., 118 U.S. 394, 396. U.S. Supreme Court decision (1949). Wheeling Steel Corporation v. Glander, 337 U.S. 562.